What are your odds of getting 80% (120 out of 150) by chance?
Answer by stanbon(75887)

What are your odds of getting 80% (120 out of 150) by chance?
---
Odds (for) = P(for)/P(against)
---
Your Problem:
Assuming there are 120 desirable items out of 150,
Odds for = 0.8/0.2 = 4:1
============================
Cheers,
Stan H.
==========